Binding energy of electrons in a metal is 250kJ/mol. Then threshold frequency of metal is?

Grade:11

4 Answers

Khushi goyal
6 years ago
electron binding energ also called ionisation potential is the energy required to remove an electron from an atom ,a molecule or an ion/. accordingly ou can work out threshold frequency by using E=hv  wher E is the bindinjg energy.
I hope this clear to you. PLEASE HIT THE APPROVED BUTTON THEN.
 
Arun
6 years ago
binding energy of electron = 250KJ/mol
binding energy of 1 electron = 250 * 103/6.023 * 1023  J
 = 4.15 * 10 –19 J
binding energy = hv0
vis thresold frequency
4.15 * 10 –19  =  6.625 * 10 –34 * v0
v0 = 6 * 1014  sec –1
Shailendra Kumar Sharma
6 years ago
Binding energy of electron =250kJ/mol
so Binding energy per electron =250*103/(6.023X1023)
=4.15X10^-19
Thus the h\nu =4.15X10-19 
thus \nu =6.27 X 1014 sec-1
RAHUL GURJAR
6 years ago
answere
binding energy of electron = 250KJ/mol
binding energy of 1 electron = 250 * 103/6.023 * 1023  J
 = 4.15 * 10 –19 J
binding energy = hv0
vis thresold frequency
4.15 * 10 –19  =  6.625 * 10 –34 * v0
v0 = 6 * 1014  sec –1